OpenStudy (anonymous):

A brick is dropped from a height of 80.0 m. a. How long does it take a brick to reach the ground?

OpenStudy (anonymous):

\[\Delta y = v _{i}\Delta t + \frac{ 1 }{ 2 }a \Delta t ^{2}\] This equation should help (assuming it is dropped from rest). You can solve for time since you have all the other variables.

OpenStudy (shamim):

The above equation is ok

OpenStudy (shamim):

U can try my equation too h=ut+(1/2)*g*t^2 Given h=80m u=0 g=9.8m/sec^2 t=?